OpenStudy (anonymous):

What is the solution to the equation log(4)x–log(4)2 = 2

OpenStudy (anonymous):

please thank you

OpenStudy (anonymous):

start by using \[\log(a)-\log(b)=\log(\frac{a}{b})\] to write \[\log_4(\frac{x}{2})=2\] then write in equivalent logarithmic form as \[\frac{x}{2}=4^2\] then solve for \(x\)

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

how do i know which #s to put where?

OpenStudy (anonymous):

not sure what you mean

OpenStudy (anonymous):

\[\large \log_b(z)=y\iff z=b^y\] you have \[\large \log_4(\frac{x}{2})=2\] which means \[\frac{x}{2}=4^2\]

OpenStudy (anonymous):

so it's 32?